OpenStudy (anonymous):

A ramp is extended from a truck to the ground, and has a slope of 7/16. If y is 3.5, what is x, in feet? 8 12.5 20 24.5 32

OpenStudy (australopithecus):

Just use the equation of the line: y = mx + b b = y-intercept (which we will set to 0) m = slope

OpenStudy (anonymous):

so it's 3/5= 7/16x+0?

OpenStudy (australopithecus):

we set the y-intercept = 0 because we assume that the latter starts at the origin of where we are measuring from

OpenStudy (australopithecus):

3.5= (7/16)x+0 solve for x

OpenStudy (anonymous):

x=8

OpenStudy (australopithecus):

ladder*

OpenStudy (anonymous):

yes

OpenStudy (australopithecus):

correct

OpenStudy (anonymous):

thnx, bro!

OpenStudy (australopithecus):

https://www.wolframalpha.com/input/?i=graph+y%3D+%287%2F16%29x